OpenStudy (anonymous):

Show that 2^99 +3^99 is divisible by 35.

OpenStudy (anonymous):

i have an inelegant method, but i am sure there is a snappier way you can factor \[x^{99}+y^{99}\] as the sum of two cubes, and continue on down when you get done one factor will be \(x+y\) which in your case is \(2+3=5\) the next factor will be \(x^2-xy+y^2\) which gives you \(7\) hello joe! now comes a better answer

OpenStudy (anonymous):

If you dont use the standard modular arithmetic way, you could write the problem as:\[(2^3)^{33}+(3^3)^{33}\]and factor it.

OpenStudy (anonymous):

yo Satellite :)

OpenStudy (anonymous):

Can you please show me how to factor the expression?

OpenStudy (anonymous):

i guess we are thinking alike nice to see you again! on break?

OpenStudy (anonymous):

\[x^n+y^n=(x+y)(x^{n-1}-x^{n-2}y+\cdots +(-1)^n y^{n-1})\]if n is odd. I might be off on the sign of the last term, but it doesnt matter for this problem.

OpenStudy (anonymous):

that gives the 5, you need the next term for the 7

OpenStudy (anonymous):

yeah, sign is wrong, that last term should be (-1)^n-1 y^n-1

OpenStudy (anonymous):

if you factor with the 33 as the major power instead of 99, you end up with the whole 35.

OpenStudy (anonymous):

maybe there is a better way. what class is this for?

OpenStudy (anonymous):

\[(2^3)^{33}+(3^3)^{33}=(2^3+3^3)(junk)\]

OpenStudy (anonymous):

much nicer

OpenStudy (anonymous):

that makes perfect sense, i did a similar question before but with difference of powers instead of the sum. Now I know that this method works for both situation. Thanks! This is math 9 and i am doing factoring difference of squares and cubes.

OpenStudy (sirm3d):

\[2^{99}+3^{99}=(2^3)^{33}+(3^3)^{33}=8^{33}+27^{33}=(8+27)(8^{32}-8^{31}27+\cdots+27^{32})\]

OpenStudy (anonymous):

yep im on break sate, and i needed it too >.> last semester was brutal =/ only one more semester till I get my bachelors though :) im applying to grad schools now, its a pain >.<

OpenStudy (anonymous):

Thank you so much for helping!!

OpenStudy (anonymous):

they should just give you the masters now then you can spend the next 7 years on phd